OpenStudy (anonymous):

PLEASE HELP! Use the following quadratic equation to complete parts a - c below. 3x2 + 2x - 5 = 0. a. Solve by factoring. Be sure to show your factored expression and then use the factors to solve for x. b. Solve using the quadratic formula. c. Describe the relationship between your answers for parts a and b.

OpenStudy (anonymous):

try \[(x-1)(3x+5)=0\] for the first one

OpenStudy (anonymous):

so \[x-1=1\implies x=1\] or \[3x+5=0\implies 3x=-5\implies x=-\frac{5}{3}\]

OpenStudy (anonymous):

thanks!